# Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits

### v1

@@ Appendix B Sketch of the Proof of Theorem 2
The proof of Theorem 2 decomposes into two steps. First, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below. Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid. Second, we show that upper and lower bounds for the $V_{t}$-norm can be transferred into upper and lower bounds for the $V_{t}^{-1}$-norm. Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound. We postpone the detailed proof for all lemmas in this section to Appendix C.
Step 1. Bounding the $V_{t}$-norm of actions.
We illustrate our geometric approach in $\mathbb{R}^{2}$ in Figure 3. We first upper bound the $V_{t}$ norm for optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$. In the following, we define $\phi_{t}=\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}-\beta_{t}^{2}}$.
@@
The proof is based on the observation that for any $\theta\in\mathcal{E}_{t}$, the line $\{t\theta:t\in\mathbb{R}\}$ (orange dashed lines in Figure 3) and ellipsoid $\mathcal{E}_{t}$ (in green) must have at least one intersection point.
In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$ (blue dashed curve). We can check this intersection is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$. A lower bound for the $V_{t}$-norm is given by the following lemma.
@@
For any $\theta\in\mathcal{H}_{t}$, we have
@@
The proof is based on the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$. Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
Figure 3: Illustration of Step 1 in $\mathbb{R}^{2}$.
Step 2. Bounding the $V_{t}^{-1}$-norm of actions
The following lemma bounds the $V_{t}^{-1}$-norm of an action $x$ given a range of its $V_{t}$-norm.
@@
Suppose $V$ has distinct eigenvalues $\lambda_{1}(V)>\lambda_{2}(V)\dots>\lambda_{d}(V)$ and $x\in\mathcal{B}_{d}$ satisfies $m\leq\|x\|_{V}^{2}\leq M$, where $M\geq 0$ is such that $\lambda_{k}(V)\leq M<\lambda_{k+1}(V)$. We have
$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$ (8)
The proof is based on writing $x$ as a linear combination of the eigenvectors of $V_{t}$, $x=\sum_{i=1}^{d}w_{i}u_{i}$, and convert the original problem into a linear programming problem with constraints using the expressions $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
Combining the two steps, we complete the proof of Theorem 2.
Appendix C Proof of Lemmas for Theorem 2
C.1 Proof of Lemma 1
Let $\theta=tx$ where $x$ is any unit vector in $\mathbb{R}^{d}$ and $t\in\mathbb{R}$ is a scalar. Consider the equation that characterizes the intersection $\{tx:t\in\mathbb{R}\}\cap\mathcal{E}_{t}$, namely $(tx-\widehat{\theta}_{t})^{\top}V_{t}(tx-\widehat{\theta}_{t})=\beta_{t}^{2}$. Equivalently, we have $t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=0$. This quadratic equation of $t$ has at least one root if the discriminant is non-negative, i.e. $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$. Then by direct computation,
Note that $\theta\in\mathcal{E}_{t}$ if and only if $\theta=tx$ for some $x$ that satisfies $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, we conclude the proof. $\square$
C.2 Proof of Lemma 2
Note that for any $\theta\in\mathcal{H}_{t}\subset\mathcal{E}_{t}$, it holds that $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$. Also, by the construction of $\mathcal{H}_{t}$, we have $\|\theta\|_{V_{t}}=\phi_{t}$. Then we conclude the proof by the definition
$\square$
C.3 Proof of Lemma 3
Let $V=U\Lambda U^{\top}$ be the eigendecomposition of $V$, so that the columns of $U$, $\{u_{1},\dots,u_{d}\}$, form an orthonormal basis of $\mathbb{R}^{d}$ and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$ is a diagonal matrix. We can write $x$ as a linear combination of $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$: $x=\sum_{i=1}^{d}w_{i}u_{i}$, where $\sum_{i=1}^{d}w_{i}^{2}=1$. By direct computation, we have $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
We would like to study the range of $\|x\|_{V^{-1}}^{2}$ when $\|x\|_{V}^{2}$ is bounded as $m\leq\|x\|_{V}^{2}\leq M$. First, let’s focus on maximizing $\|x\|_{V^{-1}}^{2}$, it suffices to solve the LP problem
maximize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}\quad\text{s.t.}\quad\forall i,a_{i}\geq 0,\ \sum_{i=1}^{d}a_{i}=1,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\leq M.$
The Lagrangian is given by
$L(d,\mu,\eta)=\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}+\mu(1-\sum_{i=1}^{d}a_{i})+\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)+\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})+\sum_{i=1}^{d}\kappa_{i}a_{i}.$
The KKT conditions are given by
$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{d}a_{i}=1,\sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{d}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$
Since $\lambda_{i}$’s are distinct, to satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices. Hence, $a_{i}$ can only be non-zero for at most two indices, denoted by $a_{i}$ and $a_{j}$. We write $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$ and hence $\lambda_{i}>\lambda_{j}$. Let $y=\|x\|_{V}^{2}\in[\lambda_{j},\lambda_{i}]$, we have $a\lambda_{i}+(1-a)\lambda_{j}=y$ and $a\lambda_{i}^{-1}+(1-a)\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
By direct computation, $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$, and clearly,
Then the maximum of $\|x\|_{V^{-1}}^{2}$ is obtained by setting $\lambda_{i}=\lambda_{1}$, $\lambda_{j}=\lambda_{d}$, $y=m$, and we have
$x^{*}_{\max}$ $=\frac{m-\lambda_{d}}{\lambda_{1}-\lambda_{d}}\lambda_{1}+\frac{\lambda_{1}-m}{\lambda_{1}-\lambda_{d}}\lambda_{d},\quad\|x^{*}_{\max}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$
Similarly, by investigating the KKT conditions for the LP problem that minimize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}$, the minimum of $\|x\|_{V^{-1}}^{2}$ is obtained by some $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$, $\lambda_{i}>\lambda_{j}$. Setting $\lambda_{u}=\min\{\lambda_{i}\geq M\}$, $\lambda_{l}=\max\{\lambda_{i}\leq M\}$, $y=M$, and we have
$x^{*}_{\min}$ $=\frac{M-\lambda_{l}}{\lambda_{u}-\lambda_{l}}\lambda_{u}+\frac{\lambda_{u}-M}{\lambda_{u}-\lambda_{l}}\lambda_{l},\quad\|x^{*}_{\min}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{u}}+\frac{1}{\lambda_{l}}-\frac{M}{\lambda_{u}\lambda_{l}}.$
This concludes the proof. $\square$
Appendix D Other Proofs
D.1 Proof of Proposition 3
By Propostion 1, we have
$\geq 1-\frac{\delta}{2}.$
Similarly, by Definition 2, we have
$\geq 1-\frac{\delta}{2}.$
Combining the two inequalities above, we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta$$\square$
D.2 Proof of Proposition 4
In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. By direct computation
$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}),$
$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}).$
Combing all terms above, we have
$\square$
D.3 Proof of Theorem 1
We formally prove Theorem 1 for completeness. The proof techniques are developed in previous papers [Abbasi-Yadkori et al., 2011, Agrawal and Goyal, 2013, Abeille et al., 2017].
This concludes the proof.$\square$
D.4 Proof of Proposition 5
This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3. Without loss of generality, we let $M>\lambda_{2}$. Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$$\square$
D.5 Proof of Theorem 3
Define $x^{\dagger}(\theta)\coloneqq\theta/\|\theta\|$. The conclusion of Theorem 2 holds for $x^{\dagger}(\theta)$, i.e. for all $t\in[1,T]$ and all $s\leq t$, we have $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s}$ and $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}$.
For any $\theta$, the feasible optimal action $x^{\star}_{s}(\theta)$ is the point in $\mathcal{X}_{s}$ that has the smallest distance from $x^{\dagger}(\theta)$. Then we have $\|x^{\dagger}(\theta)-x^{\star}_{s}(\theta)\|\leq\rho_{s}$ by the definition of the Euclidean covering radius. Therefore, let $\theta_{0}$ denote the $\theta$ that attains $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}$, it holds that
$\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}=\|x^{\star}_{s}(\theta_{0})\|_{V_{s}^{-1}}$ $\geq\|x^{\dagger}(\theta_{0})\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}).$
Similarly, we have $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$. As a result, it holds that $\alpha_{s}\leq(\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))/(\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))$$\square$
Appendix E TS-MR and Greedy-MR Algorithms

### v3

@@ 5.3 Sketch of the Proof
First note that, when $\|\widehat{\theta}_{t}\|_{V_{t}}<\beta_{t}$, the bound in Theorem 2 becomes
$\alpha_{t}\leq\lambda_{d}^{-\frac{1}{2}}(V_{t})/\lambda_{1}^{-\frac{1}{2}}(V_{t})=\sqrt{\lambda_{1}(V_{t})/\lambda_{d}(V_{t})}.$
This bound holds trivially using the fact that $\lambda_{1}(V_{t})\leq\|x\|^{2}_{V_{t}}\leq\lambda_{d}(V_{t})$. This is the case when the data is insufficient and the confidence interval is too large to get a non-trivial upper bound for $\alpha_{t}$.
In the following, without loss of generality we assume $\|\widehat{\theta}_{t}\|_{V_{t}}^{2}\geq\beta_{t}^{2}$. The proof decomposes into three steps. In the first two steps, as is illustrated in $\mathbb{R}^{2}$ in Figure 2, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below. Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid. Next, we show that upper and lower bounds for the $V_{t}$-norm can be converted into upper and lower bounds for the $V_{t}^{-1}$-norm by solving a linear programming problem. Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound. We sketch the proof below and postpone the detailed proof for all lemmas in this section to Appendix 9.
Figure 2: Illustration of Step 1 and 2 in $\mathbb{R}^{2}$. Orange dashed rays: rays starting from the origin might have different numbers of intersections with $\mathcal{E}_{t}$, indicating whether the corresponding action lies in the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. Blue dashed curve: the ellipsoid with fixed $V_{t}$-norm $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$. The intersection of this ellipsoid and $\mathcal{E}_{t}$ has the same projection as $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$.
In the following, we let $\phi_{t}\coloneqq\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-\beta_{t}^{2}}$, which is well-defined since we assume $\|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta_{t}$. In geometry, one can show that for a ray starting from the origin and intersecting $\mathcal{E}_{t}$ only at one point, $\phi_{t}$ is the ${V_{t}}$-norm of the intersection point.
Step 1. Upper Bounding the $V_{t}$-norm of actions.
Our first lemma investigates such intersection and provides an upper bound for the $V_{t}$-norm for the optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$. The proof is based on investigating the condition for a ray paralleling an action $x\in{\mathcal{S}}_{d-1}$ to intersect with $\mathcal{E}_{t}$, which means $x$ is in the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$ and might become the optimal action $x^{\star}(\theta)$.
@@
Step 2. Lower Bounding the $V_{t}$-norm of actions.
In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$. $\mathcal{H}_{t}$ consists of $\theta\in\mathcal{E}_{t}$ whose $V_{t}$-norm is $\phi_{t}$. One can check $\mathcal{H}_{t}$ is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$, and the projection of $\mathcal{H}_{t}$ onto ${\mathcal{S}}_{d-1}$ is the same as that of $\mathcal{E}_{t}$ by convexity. Hence, it suffices to only consider $\theta\in\mathcal{H}_{t}$ as the corresponding set of optimal actions coincides. A lower bound for the $V_{t}$-norm is given by the following lemma.
@@
For any $\theta\in\mathcal{E}_{t}$, we have
@@
The proof is directly using the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$. Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
Step 3. Bounding the $V_{t}^{-1}$-norm of actions
The following lemma determines the range of action $x$’s $V_{t}^{-1}$-norm based on its $V_{t}$-norm range. It turns out that the two ranges can be related using the spectral information of the sample covariance matrix $V_{t}$, which is related to the shape of the confidence ellipsoid.
@@
Let $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{V}}\}$ be the set of distinct eigenvalues of $V$ such that $\lambda_{1}>\lambda_{2}>\dots>\lambda_{N_{V}}>0$. Let $x\in\mathcal{B}_{d}$ satisfies $0<m\leq\|x\|_{V}^{2}\leq M$. We have
$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}},$ (8)
where $k$ is such that $\lambda_{k}=\max_{i\in[N_{v}]}\{\lambda_{i}\geq M\}$.
The proof involves expressing the $V$- and $V^{-1}$-norms as weighted sums of $V$’s eigenvalues, then solving a linear programming (LP) problem constrained by the norm ranges.
By inserting the upper and lower bounds of the $V_{t}$-norm from Lemmas 1 and 2 into Lemma 3, we finalize the proof of Theorem 2.
6 A Meta-Algorithm for Course-Correction
In this section, we show how the data-driven regret bound can be utilized to improve standard bandit algorithms. Specifically, we propose a meta-algorithm that constructs course-corrected variants of the base algorithms. This ensures they achieve a minimax-optimal frequentist regret guarantee (up to constant factors) while retaining the majority of the original characteristics of the base algorithms, such as computational efficiency and low regret in most cases.
We take LinTS as an example and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR). The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions when necessary.
Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Algorithm 2.